OpenStudy (anonymous):

What is 134* 6^3 / 12. Rounded?

OpenStudy (anonymous):

first you have to find 6^3 which is 6x6x6

OpenStudy (anonymous):

216

OpenStudy (jojo4eva):

rounded to what??

OpenStudy (anonymous):

yep then you multiply that by 134 x 216

OpenStudy (jojo4eva):

because the answer to the problem would be 2412

OpenStudy (jojo4eva):

what is it supposed to be rounded to. The nearest hundred? Thousand?

OpenStudy (anonymous):

yeah @jojo4eva But he was supposed to find that out himself

OpenStudy (anonymous):

yeah @jojo4eva But he was supposed to find that out himself

OpenStudy (jojo4eva):

Did it say how or what you were suppose to round too??

OpenStudy (igreen):

Is it \(\dfrac{134 \cdot 6^3}{12}\) or is it \(134 \cdot \dfrac{6^3}{12}\)

OpenStudy (anonymous):

no it did not @jojo4eva

OpenStudy (anonymous):

What are we rounding to?

OpenStudy (jojo4eva):

Then whats the point in it saying to round when its a whole number?

OpenStudy (anonymous):

It would be 2412...I think

OpenStudy (anonymous):

thats correct @Jennifer339

OpenStudy (anonymous):

:)